# A Trade-Off between Path Entanglement and Quantum Sensitivity

### v1

@@ Abstract
Introduction. Entanglement is a radically nonclassical aspect of quantum mechanics [1, 2]. It is central to our modern understanding of theoretical phenomena such as the Unruh effect and black hole radiation [3, 4, 5, 6, 7, 8], and its presence is critical for a diverse set of important applications, including quantum computation [9], teleportation [10], and imaging [11]. Its ubiquitous uses throughout quantum mechanics have led to attempts to quantify entanglement as a resource [12, 13, 14]. One such quantifier is entanglement entropy, useful for pure bipartite systems as it can be straightforwardly calculated from the system’s density matrix [15, 16, 17].
@@
We focus on entanglement in the context of measuring small phase shifts on zero-mean, multimode Gaussian states. Gaussian states are of interest experimentally, with many feasible schemes for their physical creation [23, 24, 25, 26], and also theoretically as a model for various continuous quantum phenomena [15, 17, 27]. In the setting of quantum metrology, interferometers often employ Gaussian states. Interferometers form the basis of precision measurements of gravitational waves, local gravity, and direct imaging of black holes, among others [28, 29, 30, 31]. In a typical interferometer, two input Gaussian states are first combined on a beam splitter and then propagate for a distance before recombining on a second beam splitter. The recombined states are then measured. Between the device’s beam splitters, the two paths correspond to distinct spatial modes, and each experiences a phase shift. The phase shifts are called a “common phase” if they have the same magnitude and sign and a “differential phase” if they have the same magnitude and opposite sign. When one interferometer input is a coherent state, entanglement in the form of squeezing increases its sensitivity to differential phase shifts [32].
@@
We show that in the zero-mean case, the opposite is true for entanglement between modes, i.e. between paths. Specifically, a state’s quantum Fisher information (QFI) quantifies its ultimate sensitivity to a parametrized unitary, such as a phase shift [33, 34, 35, 36]. We show that the QFI of a two-mode Gaussian state varies inversely with the entanglement entropy between the two modes. Moreover, the differential phase QFI is hurt the most by path entanglement, while the common phase QFI is unique in escaping the trade-off, due to the symmetry of the common-phase unitary. An illustration of this trade-off is that the maximally path entangled Einstein-Podolsky-Rosen (EPR) state [15] has no sensitivity at all to differential phase shifts, which is evident in the lack of active path length control in superdense coding experiments [37].
@@
We distinguish this work from that of refs. [38, 39], which find that entanglement does not improve QFI by analyzing certain specific non-Gaussian states. However, the unentangled and entangled states they consider have different numbers of modes, meaning they exist on different Hilbert spaces and transform under different unitaries. Moreover, their results are not quantified through entanglement entropy. This makes a true comparison untenable. Importantly, we arrive at a robust notion of trade-off between path entanglement and sensitivity by comparing Gaussian states on the same Hilbert space transforming under the same unitary.
@@
The covariance matrix $\mathbf{V}$ of a zero-mean state is defined as the symmetric matrix $V_{ij}:=\langle\{\hat{q}_{i},\hat{q}_{j}\}\rangle/2$. $V_{ii}$ is the variance of $\hat{q}_{i}$, and $V_{ij}$ is the covariance between $\hat{q}_{i}$ and $\hat{q}_{j}$. In this paper, we focus on zero-mean Gaussian states, which are completely characterized by their covariance matrix $\mathbf{V}$ [17, 40, 41]. This is equivalent to how multivariate Gaussian probability distributions are completely characterized by their means and covariances. In fact, as shown in Appendix B, a pure, zero-mean Gaussian state is uniquely specified among all quantum states by its covariance matrix and the zero-mean condition.
Our input state consists of a collection of independent 1-mode squeezed states with squeezing levels $r_{1},\dots,r_{N}$, characterized by the covariance matrix
@@
As discussed in Appendix B, any pure, zero-mean Gaussian state can be generated by acting on a covariance matrix of the form of $\mathbf{V}_{\text{in}}$ by a passive transformation, i.e. a photon-number preserving transformation that takes Gaussian states to Gaussian states. Each passive transformation is represented by a symplectic orthogonal matrix $\mathbf{K}$. Acting on the input state represented by $\mathbf{V}_{\text{in}}$ with this transformation yields a new state with covariance matrix $\mathbf{V}$ satisfying
@@
We are interested in the sensitivity of measuring small phase shifts on a zero-mean Gaussian state with covariance matrix $\mathbf{V}$. Specifically, assume each mode $a$ receives a phase shift of magnitude $g_{a}\phi$, and we want to measure $\phi$. For instance, we can take $g_{1}=-g_{2}=1$ to represent a differential phase shift between modes $1$ and $2$.
The state’s sensitivity to $\phi$ can be found through the quantum Fisher information (QFI), which describes how fast the state changes with respect to $\phi$ [36]. Assume that $\phi$ is tightly centered around 00, and let the QFI be $H$. The quantum Cramer-Rao bound states that the variance of a locally unbiased estimator $\phi_{\text{est}}$ of $\phi$ based on a single POVM measurement, the most general type of measurement allowed by quantum mechanics [16], must satisfy $\operatorname{Var}\phi_{\text{est}}\geq 1/\sqrt{H}$ [33, 34, 35, 36] (in the Bayesian framework, the LHS can also be turned into the mean-squared error in our estimate of $\phi$ through the van Trees inequality [44]). The upshot is that higher QFIs yield better sensitivity. As we derive in Appendix A using properties of symplectic matrices, the QFI of a pure, zero-mean Gaussian state rotating under the phase shifts $g_{a}\phi$ is
@@
For mixed states, the equality becomes an upper bound, meaning $H\leq\text{Tr}\left((\mathbf{VG})^{2}-\mathbf{G}^{2}\right)/2$. We will argue that entanglement between modes generically decreases the QFI.
Figure 1: The tradeoff between entanglement and QFI. (a) a collection of squeezed states are combined by an arbitrary multimode passive transformation $\mathbf{K}$, which in general entangles the modes. One mode then receives a phase shift $\phi$ and all modes are measured with some POVM. (b) A pair of modes are combined by an arbitrary multimode passive transformation $\mathbf{K}$, which generally entangles the modes. The first mode receives a phase shift $g_{1}\phi$, and the second receives a phase shift $g_{2}\phi$. (c) plots the scenario in (b) where the passive transformation corresponds to phase shifting one of the modes by $\theta$ and then mixing the two modes on a 50/50 beam splitter. $\theta$ controls the entanglement. As the entanglement increases, the common-phase QFI is unaffected, while the single and differential-phase QFIs decrease. This plot takes $r_{1}=r_{2}=1/2$.
Tradeoff in The Single Phase Shift Scenario. A challenge in formulating the trade-off between QFI and path entanglement for multimode systems is that multipartite entanglement is difficult to define [45]. The situation simplifies when there is a natural way to partition the system’s modes into two sets.
@@
where $\|\cdot\|^{2}$ denotes the Frobenius norm squared, the sum of the squares of the entries ([46], pg. 7). In particular,
@@
The trace and the determinant can be given a physical meaning as follows. By definition, the mean photon number of the phase-shifted mode is $n_{1}=\langle\hat{x}^{2}_{1}+\hat{p}_{1}^{2}\rangle/4-1/2$, which is related to the trace of this mode’s covariance matrix as $\operatorname{Tr}(\mathbf{V}_{1})=4n_{1}+2$.
On the other hand, the determinant is related to the entanglement between the phase-shifted mode and the rest of the system. Let $\rho_{1}$ be the reduced state achieved from tracing out all but the phase-shifted mode, and $\mu=\operatorname{Tr}(\rho_{1}^{2})$ its purity. The purity of this one-mode state monotonically decreases with its Von Neumann entropy, $S$, the standard measure of entanglement of a bipartite system. This follows from $S=\frac{1-\mu}{2\mu}\ln\left(\frac{1+\mu}{1-\mu}\right)-\ln\left(\frac{2\mu}{1+\mu}\right)$ [41]. Furthermore, the purity is related to the determinant as $\text{det}(\mathbf{V}_{1})=1/\mu^{2}$ [41] [^1] The formula in the citation differs by factors of 1/2 because we use =2Planck-constant-over-$\hbar=2$ instead of their $\hbar=1$ convention..
@@
where $H_{\text{sqz}}(n):=8n^{2}+8n$ is the QFI of a single squeezed state of photon number $n$ receiving a phase shift of $\phi$ (See Theorem 1 or [35, 48]). Now, $11/\mu^{2}-$ 1 monotonically increases with the entanglement between the phase-shifted mode and the rest of the system, yielding the trade-off between entanglement and sensitivity.
Since $\mu\leq 1$, we also get $H\leq H_{\text{sqz}}\left(n_{1}\right)$. Equality holds only when the phase-shifted mode is completely unentangled with the rest of the system, meaning $\mu=1$. Then $H$ reduces to the QFI of a 1-mode squeezed state. This makes sense because, without entanglement, the rest of the system is entirely unaffected by the phase shift, so the situation reduces to a 1-mode setup. Then eq. 7 is saying that the larger the entanglement, the further we get from this optimum.
The upper bound continues to hold for mixed states, since LABEL:QFI_general_exact is still an upper bound for mixed states. For pure states, we also have the following lower bound.
@@
The inequality follows from the fact that $\det(\mathbf{V}_{1})\leq||\mathbf{V}_{1}||^{2}/2$, which can be shown using the Cauchy-Schwarz inequality. Note that this inequality is saturated exactly when $\mathbf{V}_{1}$ is proportional to the identity, or in other words, when it represents a thermal state. Hence, this lower bound is achieved when the phase-shifted mode is maximally entangled with the rest of the system, such as for an EPR state.
@@
Tradeoff in the Two-Mode Scenario. Another system for which there is a natural bipartition is the two-mode system, as shown in fig. 1(b). In the previous section, we restricted ourselves to a phase shift on a single mode but allowed an arbitrary number of modes in the whole system. Here, we will restrict ourselves to two modes in the system but allow a phase shift on both modes.
With two modes, LABEL:QFI_general_exact yields (see Appendix A)
@@
where we have defined $H_{a}$ for $a\in\{1,2\}$ to be the QFI from a phase shift on just the $a^{\text{th}}$ mode, i.e. $g_{a}=1$ and $g_{b\neq a}=0$. Similarly, $H_{\text{com}}$ is the QFI from a common phase shift with $g_{1}=g_{2}=1$. $H_{1}$ and $H_{2}$ obey eq. 7 and are therefore affected by the entanglement between the two modes.
On the other hand, using LABEL:QFI_general_exact, $H_{\text{com}}=\operatorname{Tr}(\mathbf{V}^{2})/2-2$. This is independent of entanglement: using eq. 3, we have $\operatorname{Tr}(\mathbf{V}_{\text{in}}^{2})=\operatorname{Tr}(\mathbf{V}^{2})$, so $H_{\text{com}}$ depends only on the initial squeezed states and not the entangling transformation $\mathbf{K}$. Moreover, we can compute $H_{\text{com}}$ of the initial input states by adding their QFIs, since the inputs are independent  [^2] This is a general fact (see eq.72 in [33]) that can be seen in our case from LABEL:QFI_general_exact: $H_{\text{com}}=H_{\text{sqz}}\left(n_{\text{in},1}\right)+H_{\text{sqz}}\left(n_{\text{in},2}\right)$. In summary, eq. 10 decomposes the QFI into the sum of three QFIs based on simpler scenarios.
We can make the trade-off between QFI and entanglement clearer by substituting eq. 7 into eq. 10 and using that the purity $\mu$ is the same regardless of which mode is traced out, by the Schmidt decomposition  [^3] A standard proof of the Schmidt decomposition from pg. 704-706 in [54] can be extended to an infinite-dimensional setting with a slight addition. The addition is to note that density operators are trace class and therefore compact ([55] Theorem VI.21), so a countable spectral decomposition can still be furnished for the reduced density operators ([55] Theorem VI.16). This yields
@@
where $n_{\text{in},a}$ is the initial mean photon number in mode $a$ and $n_{\text{rot},a}$ is the mean number of photons in mode $a$ right before the phase shifts. Recall that $11/\mu^{2}-$ 1 represents the entanglement between the two modes. Therefore, the sensitivity of a differential phase measurement, where $g_{1}=-g_{2}$, is most damaged by entanglement. The common phase measurement, $g_{1}=g_{2}$, is unique in escaping the trade-off between entanglement and sensitivity.
We also note that if the initial squeezed states are equally squeezed and hence have the same photon number, after the entangling transformation $\mathbf{K}$, their photon numbers will remain the same (see proposition 3 in section C.2). In other words, we can write $n_{\text{in},1}=n_{\text{rot},1}=n_{\text{in},2}=n_{\text{rot},2}:=n$, and eq. 11 simplifies to
@@
Additionally, for the maximally path entangled EPR state, the lower bound from eq. 9 yields $21/\mu^{2}-1=H_{\text{sqz}}(n)/$ 2, so eq. 12 simplifies further to
@@
As a result, we see that EPR states are sensitive to common phase shifts but completely insensitive to differential phase shifts.
Figure 1(c) plots the sensitivity of such a pair of equally squeezed states to single, common, and differential phase shifts as their entanglement is varied. As expected, the common phase shift sensitivity is independent of entanglement. Meanwhile, when the state is maximally path entangled, the single mode phase shift drops to half its maximal value while the differential phase sensitivity vanishes.
Decoupled Squeezed States Optimize the QFI. In the $N$-mode case with general $g_{a}$ values, there are many relevant notions of path entanglement because there is no clear bipartition, making it difficult to write a formula for the QFI as a function of the path entanglement. However, there is always a well-defined notion of no entanglement: the state decomposes as a tensor product  [^4] In cases with identical particles, the definition of no entanglement can sometimes be more subtle due to the need to (anti-)symmetrize [56]. However, since we are employing a second quantization approach, these symmetrization issues do not arise. In particular, we are interested in entanglement between (spatial) modes, and these modes are concatenated straightforwardly through the tensor product.. We show that the QFI attains its maximum when path entanglement is absent.
To formalize this, we say $\mathbf{V}$ is decoupled when the covariances vanish between all pairs of modes. In our setup, this condition is equivalent to the state having no path entanglement:
@@
For a proof, see Appendix C.1.
We say that a decoupled state is properly ordered if given any modes $a$ and $b$, if $g_{a}^{2}>g_{b}^{2}$, then mode $a$ is at least as squeezed as mode $b$. For convenience, let’s assume the input squeezing levels satisfy $r_{1}\geq\dots\geq r_{N}$ (since we can always permute the input states using $\mathbf{K}$), and label the modes such that the rotation rates satisfy $g_{1}^{2}\geq\ldots\geq g_{N}^{2}$. We have (see appendix C for the proof)
@@
For instance, since $\mathbf{V}_{\text{in}}$ is decoupled and properly ordered, if we do not act with a passive transformation (i.e. $\mathbf{K}=I$), we will have an optimal QFI state.
As a consequence of Theorem 1, the QFI of a pure, zero-mean Gaussian state passively transformed from 1-mode squeezed states with squeezing of at most $r$ is bounded by
@@
This bound demarcates the maximal phase sensitivity achievable with any passive symplectic transformation based on the initial collection of squeezed states.
@@
We also note that eq. 15 generally will not hold if we remove the zero-mean condition. This is because the maximal squeezing level present in the initial squeezed states cannot be increased by a passive transformation, whereas if the initial states were displaced, passive transformations could increase their maximal displacement. As a result, with enough modes, it can become advantageous to inject displaced states instead of squeezed vacuum states, and the bound of eq. 15 can be evaded. Indeed, if we take $g_{1}=1$ and $g_{a}=0\ \forall a>1$, the RHS of eq. 15 is independent of $N$. However, if one allows a nonzero displacement in each mode, one can use $\mathbf{K}$ to siphon all the displacement into the phase-shifted mode, completely flouting the spatial mean photon number constraint and ensuring the QFI scales as $\sqrt{N}$. Then with enough modes, the displacement strategy will become more effective.
As another corollary of Theorem 1, since a 1-mode squeezed state is trivially decoupled and properly ordered, its QFI immediately follows as $H=2g_{1}^{2}\sinh^{2}(2r)$. Since for a 1-mode squeezed state, $n=\sinh^{2}(r)$, if we take $g_{1}=1$, we get the familiar $H_{\text{sqz}}(n)=8n^{2}+8n$. In general, letting $n_{\text{in},i}$ be the mean photon number in mode $i$, we could then write the optimal QFI as

### v2

@@ Abstract
Introduction. Entanglement is a radically nonclassical aspect of quantum mechanics [1, 2]. It is central to our modern understanding of theoretical phenomena such as the Unruh effect and black hole radiation [3, 4, 5, 6, 7, 8], and its presence is critical for a diverse set of important applications, including quantum computation [9], teleportation [10], and imaging [11]. Its ubiquitous uses throughout quantum mechanics have led to attempts to quantify entanglement as a resource [12, 13, 14].
@@
We focus on entanglement in the context of measuring small phase shifts on zero-mean, multimode Gaussian states. Gaussian states are of interest experimentally, with many feasible schemes for their physical creation [20, 21, 22, 23], and also theoretically as a model for various continuous quantum phenomena [24, 25, 26]. In the setting of quantum metrology, interferometers often employ Gaussian states.
@@
Interferometers form the basis of precision measurements of gravitational waves, local gravity, and direct imaging of black holes, among others [27, 28, 29, 30]. In a standard interferometer, two input Gaussian states are first combined on a beam splitter and then propagate for a distance before recombining on a second beam splitter. The recombined states are then measured. Between the device’s beam splitters, the two paths correspond to distinct spatial modes, and each path experiences a phase shift. The phase shifts are called a “common phase” if they have the same magnitude and sign and a “differential phase” if they have the same magnitude and opposite sign. In interferometers with non-zero mean Gaussian inputs or non-Gaussian inputs, differential phase sensitivity often improves as path entanglement increases [31, 32, 33, 34, 35].
@@
We show that the opposite is true for zero-mean Gaussian states: there is a trade-off between path entanglement and optimal phase sensitivity. Specifically, the quantum Fisher information (QFI) of a state quantifies its ultimate sensitivity to a parametrized unitary such as a phase shift [36, 37, 38, 39]. We show that the QFI of a two-mode Gaussian state varies inversely with the entanglement entropy between the two modes. Moreover, the differential phase QFI is reduced the most by path entanglement, while the common phase QFI is unique in escaping the trade-off, due to the symmetry of the common phase unitary.
We then generalize these results to $N$-mode interferometers using zero-mean Gaussian states and show that path entanglement generically reduces their phase sensitivity. We determine the states that optimize QFI and find that in all cases, we can reach optimal sensitivity with zero-mean Gaussian states possessing no path entanglement.
We note that zero-mean Gaussian states are of particular interest experimentally; it has been shown that they achieve the optimal differential phase sensitivity among all initially unentangled input states with a fixed mean photon number in standard two-mode interferometers [40].
We can gain intuition for why path entanglement degrades phase sensitivity in zero-mean Gaussian states’ by comparing squeezed states with Fock states. Fock states are invariant with respect to phase shifts, but path entanglement breaks this invariance, improving their phase sensitivity. On the other hand, unentangled squeezed states naturally have their major and minor axes in a plane defined by phase shifts. Gaussian path-entangling transformations rotate the squeezed states in a higher dimensional phase space so that phase shifts no longer rotate the most anti-squeezed major axis into the most squeezed minor axis. This effectively reduces the amount of squeezing and, hence, their phase sensitivity.
@@
We distinguish this work from that of refs. [41, 42], which argue that entanglement does not improve QFI by comparing certain specific, non-Gaussian states. We arrive at a robust notion of trade-off between path entanglement and sensitivity by comparing Gaussian states on the same Hilbert space transforming under the same unitary.
@@
The covariance matrix, $\mathbf{V}$, of a zero-mean state is defined as the symmetric matrix $V_{ij}:=\langle\{\hat{q}_{i},\hat{q}_{j}\}\rangle/2$. $V_{ii}$ is the variance of $\hat{q}_{i}$, and $V_{ij}$ is the covariance between $\hat{q}_{i}$ and $\hat{q}_{j}$. In this paper, we focus on zero-mean Gaussian states, which are completely characterized by their covariance matrix $\mathbf{V}$ [25, 43, 44]. This is equivalent to how multivariate Gaussian probability distributions are completely characterized by their means and covariances. In fact, as shown in Supplementary Information section II, a pure, zero-mean Gaussian state is uniquely specified among all quantum states by its covariance matrix and the zero-mean condition.
Our input state is a collection of independent squeezed states with squeezing levels $r_{1},\dots,r_{N}$, characterized by the covariance matrix
@@
As discussed in Supplementary Information section II, any pure, zero-mean Gaussian state can be generated by acting on a covariance matrix in the form of $\mathbf{V}_{\text{in}}$ by a passive transformation, i.e. a photon-number preserving transformation that takes Gaussian states to Gaussian states. Each passive transformation is described by a symplectic orthogonal matrix $\mathbf{K}$. This transformation takes the input state, $\mathbf{V}_{\text{in}}$, to a new state with covariance matrix $\mathbf{V}$ satisfying
@@
We are interested in the sensitivity of measuring small phase shifts on a zero-mean Gaussian state with covariance matrix $\mathbf{V}$. Specifically, assume each mode $a$ receives a phase shift of magnitude $g_{a}\phi$, and we want to measure $\phi$. For instance, $g_{1}=-g_{2}=1$ represents a differential phase shift between modes $1$ and $2$. While we focus on phase shifts for concreteness, we note that similar results hold for any parametrized passive symplectic transformation (see Supplementary Information section IV).
The state’s sensitivity to a phase shift parametrized by $\phi$ can be quantified through the quantum Fisher information (QFI). The QFI describes how fast the state changes with respect to $\phi$ [39]. Specifically, fix $\phi=\phi_{0}$ and let the QFI be $H(\phi_{0})$. Let $\phi_{\text{est}}$ be an estimator of $\phi$ satisfying $\partial_{\phi}\langle\phi_{\text{est}}\rangle|_{\phi=\phi_{0}}=1$, and assume the estimator is based on a single POVM measurement, the most general type of measurement allowed by quantum mechanics [47]. The quantum Cramer-Rao bound states that the variance of $\phi_{\text{est}}$ must satisfy $\operatorname{Var}\phi_{\text{est}}\geq 1/\sqrt{H(\phi_{0})}$ [48, 37, 38, 39]. In the Bayesian framework, if $\phi$ is tightly centered around $\phi_{0}$, the LHS can also essentially be replaced with the mean-squared error in any estimate of $\phi$ through the van Trees inequality [49]. The upshot is that higher QFIs yield better sensitivity.
As derived in Supplementary Information section I using properties of symplectic matrices, the QFI of a pure, zero-mean Gaussian state rotating under the phase shifts $g_{a}\phi$ is
@@
Note that this is independent of $\phi_{0}$. For mixed states, the equality becomes an upper bound, $H\leq\text{Tr}\left((\mathbf{VG})^{2}-\mathbf{G}^{2}\right)/2$. We will argue that entanglement between modes generically decreases the QFI.
Figure 1: The trade-off between entanglement and QFI. (a) a collection of squeezed states combine through an arbitrary multimode passive transformation $\mathbf{K}$, which generally entangles the modes. One mode then receives a phase shift $\phi$. Finally, all modes are measured with some POVM. (b) a pair of modes are combined by an arbitrary multimode passive transformation $\mathbf{K}$, which generally entangles the modes. The first mode receives a phase shift $g_{1}\phi$, and the second receives a phase shift $g_{2}\phi$. (c) plots the QFI from scenario (b) where the passive transformation corresponds to phase shifting one of the modes by $\theta$ and then mixing the two modes on a 50/50 beam splitter. $\theta$ controls the entanglement. As the entanglement increases, the common-phase QFI is unaffected, while the single and differential-phase QFIs decrease. This plot takes $r_{1}=r_{2}=1/2$.
Trade-off in the Single Phase Shift Scenario. A challenge in formulating the trade-off between QFI and path entanglement for multimode systems is that multipartite entanglement is difficult to define [50]. The situation simplifies when there is a natural way to partition the system’s modes into two sets.
@@
where $\|\cdot\|^{2}$ denotes the Frobenius norm squared, the sum of the squares of the entries. In particular,
@@
The trace and determinant have physical meaning. By definition, the phase-shifted mode has mean photon number $n_{1}=\langle\hat{x}^{2}_{1}+\hat{p}_{1}^{2}\rangle/4-1/2$, so the trace of this mode’s covariance matrix is given by $\operatorname{Tr}(\mathbf{V}_{1})=4n_{1}+2$.
On the other hand, the determinant, $\det(\mathbf{V}_{1})$, is related to the entanglement between the phase-shifted mode and the rest of the system. Let $\rho_{1}$ be the reduced state achieved from tracing out all but the phase-shifted mode, and $\mu=\operatorname{Tr}(\rho_{1}^{2})$ its purity. The purity of this one-mode state monotonically decreases with its Von Neumann entropy, $S$, the standard measure of entanglement of a bipartite system. This follows from $S=\frac{1-\mu}{2\mu}\ln\left(\frac{1+\mu}{1-\mu}\right)-\ln\left(\frac{2\mu}{1+\mu}\right)$ [44]. Furthermore, the purity is related to the determinant as $\text{det}(\mathbf{V}_{1})=1/\mu^{2}$ [44] [^1] The formula in the citation differs by factors of 1/2 because we use =2Planck-constant-over-$\hbar=2$ instead of their $\hbar=1$ convention..
@@
where $H_{\text{sqz}}(n):=8n^{2}+8n$ is the QFI of a single squeezed state of photon number $n$ receiving a phase shift of $\phi$ (see above eq. 16 or [38, 52]). Now, $11/\mu^{2}-$ 1 monotonically increases with the entanglement between the phase-shifted mode and the rest of the system, yielding a trade-off between entanglement and sensitivity.
Since $\mu\leq 1$, we also find $H\leq H_{\text{sqz}}\left(n_{1}\right)$. Equality holds only when the phase-shifted mode is completely unentangled with the rest of the system, since then $\mu=1$. In this unentangled case, $H$ reduces to the QFI of a 1-mode squeezed state. This makes sense because, without entanglement, the rest of the system is entirely unaffected by the phase shift, so the situation reduces to a 1-mode setup. Then eq. 7 expresses that the larger the entanglement, the further we get from this optimum.
Our upper bound still holds for mixed states, because LABEL:QFI_general_exact remains an upper bound. For pure states, we also have the following lower bound
@@
The inequality follows from the fact that $\det(\mathbf{V}_{1})\leq||\mathbf{V}_{1}||^{2}/2$, which can be shown using the Cauchy-Schwarz inequality, and eq. 6. The lower bound is saturated when $\mathbf{V}_{1}$ is proportional to the identity, or in other words, when it represents a thermal state. Hence, this lower bound is achieved when the phase-shifted mode is maximally entangled with the rest of the system, such as for an EPR state.
@@
Trade-off in the Two-Mode Scenario. In the preceding section, the state had an arbitrary number of modes and was bipartite because only one mode was phase-shifted. Here, we instead consider two-mode states, as depicted in fig. 1(b). Two-mode states are naturally bipartite, so we can allow a phase shift on both modes.
With two modes, LABEL:QFI_general_exact yields (see Supplementary Information section I)
@@
where we have defined $H_{a}$ for $a\in\{1,2\}$ to be the QFI from a phase shift on just the $a^{\text{th}}$ mode, i.e. $g_{a}=1$ and $g_{b\neq a}=0$. Similarly, $H_{\text{com}}$ is the QFI from a common phase shift, where $g_{1}=g_{2}=1$. $H_{1}$ and $H_{2}$ obey eq. 7 and are therefore affected by the entanglement between the two modes.
Furthermore, LABEL:QFI_general_exact also yields $H_{\text{com}}=\operatorname{Tr}(\mathbf{V}^{2})/2-2$. This is independent of entanglement: using eq. 3, we have $\operatorname{Tr}(\mathbf{V}_{\text{in}}^{2})=\operatorname{Tr}(\mathbf{V}^{2})$, so $H_{\text{com}}$ depends only on the initial squeezed states and not the entangling transformation $\mathbf{K}$. Moreover, we can compute $H_{\text{com}}$ of the initial input states by adding their QFIs, since the inputs are independent [36]: $H_{\text{com}}=H_{\text{sqz}}\left(n_{\text{in},1}\right)+H_{\text{sqz}}\left(n_{\text{in},2}\right)$. In summary, eq. 10 decomposes the QFI into the sum of three QFIs based on simpler scenarios.
We can make the trade-off between QFI and entanglement clearer by substituting eq. 7 into eq. 10 and noting that tracing out either mode yields the same purity, $\mu$, by the Schmidt decomposition  [^2] We can extend the proof of the Schmidt decomposition from pg. 704-706 in [61] to an infinite-dimensional setting by noting that density operators are trace class and therefore compact ([62] Theorem VI.21). This means a countable spectral decomposition can still be furnished for the reduced density operators ([62] Theorem VI.16), which is a key step of the proof.. We obtain
@@
where $n_{\text{in},a}$ and $n_{\text{rot},a}$ are the mean photon numbers in mode $a$ of the initial state, $\mathbf{V}_{\text{in}}$, and the entangled state, $\mathbf{V}$, respectively. Recall $11/\mu^{2}-$ 1 is monotonic with the entanglement between the two modes. Therefore, the sensitivity of differential phase measurements, $g_{1}=-g_{2}$, is most damaged by entanglement. The common phase measurement, $g_{1}=g_{2}$, is unique in escaping the trade-off between entanglement and sensitivity. Indeed, because the common phase transformation commutes with the entangling transformation $\mathbf{K}$, path entanglement cannot affect its QFI.
The formula simplifies when the initial squeezed states are equally squeezed and thus have the same mean photon number. After the entangling transformation $\mathbf{K}$, the mean photon numbers of the two modes will remain the same (see Supplementary Information section III B). Therefore, we can write $n_{\text{in},1}=n_{\text{rot},1}=n_{\text{in},2}=n_{\text{rot},2}:=n$, simplifying eq. 11 to
@@
Additionally, for the maximally path entangled EPR state, the saturation of the lower bound from eq. 9 yields $21/\mu^{2}-1=H_{\text{sqz}}(n)/$ 2, so eq. 12 simplifies further to
@@
Thus, EPR states are sensitive to common phase shifts, where $g_{1}=g_{2}$, but completely insensitive to differential phase shifts, where $g_{1}=-g_{2}$.
Figure 1(c) plots the sensitivity of a pair of equally squeezed states to single, common, and differential phase shifts as the path entanglement is varied. As expected, the common phase sensitivity is independent of entanglement. Meanwhile, when the state is maximally path entangled, the single mode phase sensitivity is halved while the differential phase sensitivity vanishes.
Decoupled Squeezed States Optimize the QFI. In the $N$-mode case with general $g_{a}$ values, many relevant notions of path entanglement exist due to the lack of a natural bipartition, making it challenging to give a formula for the QFI as a function of path entanglement. However, there is always a well-defined notion of no entanglement: the state decomposes as a tensor product. We show that the QFI attains its maximum when path entanglement is absent.
To formalize this, we say $\mathbf{V}$ is decoupled when the covariances vanish between all pairs of modes. In our setup, this condition is equivalent to the absence of path entanglement:
@@
For a proof, see Supplementary Information section III A.
We say a state is properly ordered if for any modes $a$ and $b$, if $g_{a}^{2}>g_{b}^{2}$, then their mean photon numbers satisfy $n_{a}\geq n_{b}$. For convenience, label the modes such that $g_{1}^{2}\geq\ldots\geq g_{N}^{2}$. Furthermore, by absorbing permutations of modes into $\mathbf{K}$, we may assume the input squeezing levels satisfy $r_{1}\geq\dots\geq r_{N}$. We have (proof in Supplementary Information section III):
@@
For instance, since $\mathbf{V}_{\text{in}}$ is already decoupled and properly ordered, it is an optimal QFI state; we need not perform a passive transformation at all to reach the optimal QFI.
Theorem 1 demarcates the maximal phase sensitivity achievable with any passive transformation based on the initial collection of squeezed states. Constraining the squeezing of the initial states to at most $r$, the bound becomes
@@
We also note that eq. 15 breaks down if we remove the zero-mean condition. This is because the maximal squeezing level present in the initial squeezed states cannot be increased by a passive transformation, whereas if the initial states were displaced, passive transformations could increase their maximal displacement. As a result, with enough modes, it can become advantageous to inject displaced states instead of squeezed vacuum states, evading the bound. Indeed, taking $g_{1}=1$ and $g_{a}=0\ \forall a>1$, the RHS of eq. 15 is independent of $N$. However, with displacement in the initial states, $\mathbf{K}$ can siphon all the displacement into the phase-shifted mode, ensuring the QFI scales as $N$. Then with enough modes, the displacement strategy will become more effective.
As another corollary of Theorem 1, since a 1-mode squeezed state is trivially decoupled and properly ordered, its QFI immediately follows as $H=2g_{1}^{2}\sinh^{2}(2r)$. Since for a 1-mode squeezed state, $n=\sinh^{2}(r)$, by taking $g_{1}=1$, we find $H_{\text{sqz}}(n)=8n^{2}+8n$. In general, letting $n_{\text{in},i}$ be the mean photon number in mode $i$, we could then rewrite the optimal QFI as